# Carrer Pic de Peguera 13, 17003 Girona, Spain

### v1

@@ 3 Simulation Setup
This section describes the simulation environment, sensor and vehicle configurations, and the data collection approach for generating the event-based optical flow prediction dataset, eStonefish-scenes, tailored for autonomous underwater vehicles. The dataset consists of event streams, grayscale images, and the corresponding ground truth optical flow. Additionally, it utilizes an optimized format provided within the eWiz framework to facilitate efficient data storage, access, and processing. Further details on this optimized format are provided in Section 5.
The dataset was generated using the Stonefish simulator (Grimaldi et al., 2025) within a custom-designed environment featuring a texture-rich seabed. To add visual complexity, the seabed was populated with abundant corals and marine flora with pronounced edges and contrasts. The corals were divided into 3 clusters distributed across the seabed, with each cluster containing 20 corals of different sizes, ranging from small antler corals to large dome corals. The scene was randomly generated using a scene generator to diversify coral pose and size. The dataset features two environments:
  • Rocky: Datasets with the rocky environment capture a rocky seabed texture. It is characterized by a barren rocky seabed with no flora in the scene.
  • Reef: Datasets with the reef environment are characterized by the same rocky seabed texture, but populated with flora to simulate the look and feel of a colourful coral reef.
The dataset comprises both, static and dynamic scenes. Static scenes are simple, empty environments with no dynamic objects, such as schools of fish. Only stationary corals appear in these scenes, ensuring consistency with event-based optical flow loss functions commonly utilized in the literature (Stoffregen and Kleeman, 2019). Dynamic scenes, on the other hand, incorporate dynamic elements in the form of schools of fish. Several fish models were used across the simulation, each exhibiting slightly modified Boid behaviours (Reynolds, 1987) to mimic the behaviour of schools of fish in nature. This diversity in scene complexity, creating complex visual patterns, provides a comprehensive framework for both training and evaluating optical flow algorithms across various environments. Figure 1 depicts an example of the generated scene.
Figure 1: Generated environment, the environment includes a seabed with feature-rich corals
For data collection, a joystick-controlled BlueROV2 (Blue Robotics, 2016) model was utilized as shown in Figure 2. The ROV was equipped with an RGB camera, an event-based camera, and an optical flow sensor, mounted in two configurations. To ensure diverse and robust data, the scene was recorded from multiple viewpoints, mimicking the placement of a camera on a real ROV. Specifically, the sensors were positioned in the following configurations:
@@
  • Down-looking: In this setup, the ROV maintained a constant depth while moving, with a focus on the seabed.
The forward-looking configuration is particularly advantageous for tasks such as scene inspection and obstacle avoidance, as it provides a clear view of the environment ahead. In contrast, the down-looking configuration is optimized for computing the odometry of the robot, leveraging the EBC feed to analyse the textured seabed. These complementary configurations enable the ROV to address a diverse range of underwater exploration and navigation challenges.
@@
[…]
@@ 2 Static Varying Reef Front-looking
3 Static Constant Rocky Down-looking
@@ 4 Static Varying Rocky Front-looking
5 Dynamic Constant Reef Down-looking
@@ 6 Dynamic Varying Reef Front-looking
7 Dynamic Constant Rocky Down-looking
@@ 8 Dynamic Varying Rocky Front-looking
Table 1: Summary of the generated sequences and their corresponding configurations.
These sequences can be utilized for training neural networks by dividing them into separate event volumes. Each event volume contains a stream of events, along with all grayscale images that occurred at the corresponding timestamps. One such example is by extracting events that took place between every pair of consecutive grayscale images. The resulting volumes would include a grayscale frame at time $t$, another frame at time $t+1$, and all events that happened in between. Based on the 20 FPS refresh rate of the grayscale sensor, we can extract 800 training samples per sequence approximately, providing around 6400 samples for all sequences combined.
@@ 4 Data Generation Pipeline
The dataset can be easily extended by using a modular data generation pipeline, which comprises 3 separate packages, each responsible for different aspects of data generation. First, we present the Stonefish-scenegen package. This package is used to generate coral-rich and vibrant scenes by randomly populating the seabed with corals of different types and sizes. It generates a Stonefish-compatible scene file, which can then be used to launch the main pipeline, eStonefish-scenes. The pipeline is a ROS package with a launch file for Stonefish. The launch file starts the simulator with the generated environment, spawning the ROV and displaying live feeds from all its sensors. The pipeline internally uses the Stonefish-boids package to randomly spawn schools of fish in the scene. The user can then move the ROV around the scene while recording a rosbag, which can subsequently be converted to the eWiz format using the ConvertStonefish module of eWiz.
[…]
@@ 4.1 Scene Generation
One challenge with the Stonefish simulator is the time-consuming process of creating rich and dynamic underwater scenes. Each coral or 3D object must be manually defined in a scene file, specifying its pose in the world frame, size, and physical attributes. While this approach is feasible for a few objects, it becomes impractical for environments featuring numerous objects. To address this limitation, we introduce Stonefish-scenegen, an underwater scene generator specifically designed for the Stonefish simulator. This tool streamlines the creation of complex underwater environments, significantly reducing the effort required for scene generation. It consists of a Python script that takes as input a Stonefish scene file, a ground surface, and the desired 3D models to spawn. The user chooses the number of coral clusters, along with the number of corals spawned in each cluster. Each coral instance is placed by sampling a random vertex on the ground surface. The normal of the surface passing through that vertex is then used to determine translation and rotation transformations. To ensure reproducibility, the user can generate the exact same scene by fixing a seed before running the script. The script also converts from the ENU coordinate system to its NED equivalent, as required by Stonefish. Finally, a scene file containing the path, texture, size, and pose of all spawned 3D models is generated. This scene file can then be used with eStonefish-scenes to launch the simulation. The scene generator is made openly available at the following link: https://github.com/CIRS-Girona/stonefish-scenegen.
@@ 4.2 Simulating Schools of Fish
To enhance the visual complexity of the underwater scenes with some dynamic elements, we incorporated an adapted implementation of the Boids algorithm (Reynolds, 1987) for simulating the natural grouping behaviour observed in schools of fish. The algorithm is based on four fundamental rules that govern the behaviour of each fish within the school, collectively influencing the behaviour of the group as a whole.
@@
  • Separation: Each fish receives a repulsion force if it gets too close to another fish in the school. This helps maintain a minimum distance between all fish.
  • Leader Attraction: One fish is designated as the leader of the school and follows a predefined path along the map, while the other fish experience an attractive force between them and the leader.
@@
To efficiently handle such environmental interactions with static objects, we use OctoMap (Hornung et al., 2013) for discretizing the simulation space, as illustrated in figure 4. This enables rapid detection of potential collision zones and the assignment of repulsive force vectors to different locations. The flexibility of the library allows for the creation of a custom Octree where each node stores a repulsive force vector normal to the object surface. When a fish enters an obstacle voxel, a repulsive force is applied, ensuring smooth and natural avoidance behaviour. For dynamic objects such as the ROV, on the other hand, repulsive forces are applied based on the distance between each fish and a defined repulsion radius around the object, allowing for real-time interaction.
@@
Figure 4: Discretized underwater environment as an Octree structure.
Additionally, each school of fish follows a randomly generated path created using the Rapidly-exploring Random Trees (RRT) algorithm (LaValle, 1998), implemented through the Open Motion Planning Library (OMPL) (Şucan et al., 2012). The designated leader follows this path, while the rest of the group moves along with it based on the leader attraction rule. Each time a new simulation is initiated, both the path and the spawning positions of the fish are randomly generated. However, for reproducibility, users can set a fixed seed to ensure the path and spawning positions remain consistent across runs.
Users can customize the type of fish schools and adjust various behavioral parameters through a YAML configuration file. These include alignment, cohesion, and separation weights, as well as spawning location, model scale, and school size. This functionality is provided as an open-source ROS package, Stonefish-boids, which builds on the base Stonefish API. The package is available at: https://github.com/CIRS-Girona/stonefish_boids.
@@ 4.3 Data Collection
To record data from the generated scenes, we provide a ROS-based utility, eStonefish-scenes, which uses the scene file generated by the scene generator and internally calls stonefish-boids to simulate schools of fish. It relies on the Stonefish-ROS bridge to facilitate communication with the simulator. The package also includes meshes and textures for the BlueROV2 and a rocky seabed. A predefined scenario combines these assets into a single environment, spawning the ROV and the seabed beneath. A Python script is provided to control the robot, utilizing its allocation matrix to enable joystick-based operation. Additionally, an events rendering script is included to visualize the resulting event image in RViz (Kam et al., 2015) as the simulation progresses. Users can customize the sensors mounted on the ROV by modifying the corresponding scene file. Customization options include, but are not limited to, adjusting resolution, field of view, event-based sensor thresholds, and adding noise. The record.bash script allows users to record rosbags from the simulation, which can subsequently be converted to the eWiz format using the ConvertStonefish module of eWiz. This data collection utility is available at the following link: https://github.com/CIRS-Girona/estonefish-scenes.
@@ 5 Data Processing
The eStonefish-scenes dataset was generated with NN training and inference in mind. To streamline this process, we created eWiz, a Python-based library that can be installed with PyPi (The Python Community, 2003), serving as an all-inclusive framework for event-based data manipulation, processing, visualization, and evaluation. eWiz is designed for seamless integration with libraries such as PyTorch (The PyTorch Foundation, 2016) and Tonic (Gregor Lenz, 2019), making it suitable for ML-oriented pipelines. Additionally, it offers compatibility with existing datasets and supports both event-based and frame-based data. The following subsections present an overview of the various functionalities currently offered by eWiz. For further details, the readers are encouraged to refer to the official documentation. The library is openly available at the following link: https://github.com/CIRS-Girona/ewiz.
Figure 5: Currently implemented modules for the eWiz library.
@@ 5.1 Data Access and Storage
Due to the high temporal resolution and data throughput of EBCs, event-based data requires substantial storage space. Additionally, slicing data between any two timestamps requires time-consuming search algorithms due to the large volume of events in the sequence. Libraries like NumPy Harris et al., 2020 also face high memory demands, as they typically require loading entire sequences into memory before slicing or processing. The eWiz format addresses these challenges with an optimized approach to data management and manipulation. Firstly, we use Blosc (The Blosc Developers, 2014), a high-performance compressor optimized for binary data, effective for compressing numerical arrays. To further reduce data size, we constrain the possible values for every event stream. For instance, pixel coordinates are stored as 16-bit unsigned integers, while polarities are stored as boolean values. Next, we use the h5py library (Andrew Collette and Contributors, 2017) read and write data as HDF5 files (The HDF Group, 2006), which supports large and complex data structures, and allows slicing without the need to completely load the data into memory. Finally, to reduce loading times, we pre-compute the mappings between timestamps and data indices during the data writing process. This includes binary search to map timestamps (in $ms$) to event, grayscale, and flow indices, as well as direct mappings of event indices to their grayscale and flow counterparts. Additionally, optical flow data is automatically synchronized between specified timestamps by interpolating and accumulating flow displacements as needed.
eWiz provides easy-to-use readers and writers to facilitate this process. The data reader uses the precomputed time mappings to read the desired indices of events based on the start and end timestamps and loads them to memory. It also loads the corresponding grayscale images that were captured at the beginning and at the end of the event stream along with the synchronized optical flow. The user has the possibility to stride over the events, by either using event indices, timestamps, or grayscale images indices. The data reader goes hand in hand with the included data iterators which allow to sequentially stride over the data. Such functionality allows the user to obtain event-based data sequentially, without the need for manual indexing.
eWiz also aims to support a multitude of EBCs and ensure compatibility with publicly available datasets. We include several data converters for transforming these data types into the eWiz format. Currently, we support the DAVIS346 (Inivation, 2019) and Prophesee (Prophesee, 2024) cameras, along with popular datasets such as MVSEC (Zhu et al., 2018a) and DSEC (Gehrig et al., 2021).

### v2

@@ 3 Simulation Setup
This section describes the simulation environment, sensor and vehicle configurations, and the data collection approach for generating the eStonefish-Scenes dataset. The dataset consists of event streams, grayscale images, and the corresponding ground-truth optical flow. Additionally, it utilizes an optimized format provided within the eWiz framework to facilitate efficient data storage, access, and processing. Further details on this optimized format are provided in Section 5.
The dataset was generated using the Stonefish simulator (stonefishCielak19; stonefishGrimaldi25) with a custom-designed environment featuring a texture-rich seabed. To add visual complexity, the seabed was populated with abundant corals and marine flora featuring pronounced edges and contrasts. The corals were divided into 3 clusters distributed across the seabed, with each cluster containing 20 corals of different sizes, ranging from small branching corals to large dome corals. The scene was randomly generated using a scene generator to diversify coral poses and sizes. The dataset features two environments:
  • Rocky: Sequences characterized by a barren rocky seabed with no flora in the scene.
  • Reef: Sequences characterized by the same rocky seabed texture, but populated with flora to simulate the look and feel of a colourful coral reef.
The dataset comprises both static and dynamic scenes. Static scenes are simple environments with no dynamic objects, such as schools of fish. Only stationary corals appear in these scenes, ensuring consistency with event-based optical flow loss functions commonly utilized in the literature (mcStoffregen19). Dynamic scenes, on the other hand, incorporate moving elements in the form of schools of fish. Several fish models were used across the simulation, each exhibiting slightly modified Boid behaviours (boidsReynolds87) to mimic natural schooling behaviour of fish in nature. This diversity in scene complexity, which creates complex visual patterns, provides a comprehensive framework for both training and evaluating optical flow algorithms across various environments. Figure 1 depicts an example of the generated scene.
Figure 1: Generated environment: the environment includes a seabed with feature-rich corals.
For data collection, a joystick-controlled BlueROV2 (bluerov2) model was utilized, as shown in Figure 2. The ROV was equipped with an RGB camera, an event-based camera, and an optical flow sensor, all of which are mounted in two configurations. Unlike the RGB and EBC cameras, that are modelled after real sensors in Stonefish, the optical flow sensor is a pseudo-sensor that uses the information from the 3D environment to create a perfect, noise-free representation of the optical flow data. To ensure diverse and robust data, the scene was recorded from multiple viewpoints, mimicking the placement of a camera on a real ROV. Specifically, the sensors were positioned in the following configurations:
@@
  • Down-looking: In this setup, the ROV maintained a constant depth while moving, focusing on the seabed.
The forward-looking configuration is particularly advantageous for tasks such as scene inspection and obstacle avoidance, as it provides a clear view of the environment ahead. In contrast, the down-looking configuration is optimized for computing the robot odometry, leveraging the EBC feed to analyse the textured seabed. These complementary configurations enable the ROV to address a diverse range of underwater exploration and navigation challenges.
@@
All cameras mounted on the ROV have a resolution of $346\times 260$ pixels and a field-of-view of $70\text{\,}\mathrm{\SIUnitSymbolDegree}$. The RGB, EBC, and optical flow sensors share the same intrinsic and extrinsic camera parameters, and can be considered as a single camera that produces perfectly aligned colour, event, and optical flow data. The EBC was configured with positive and negative thresholds of 0.28, and a refractory period of $200\text{\,}\mathrm{ns}$. The simulation clock was set to $300\text{\,}\mathrm{Hz}$. An NVIDIA RTX 4090 GPU was used for the simulation and data collection. This configuration capped the grayscale and optical flow sensors at $20\text{\,}\mathrm{Hz}$, while the event-based camera compared the frame brightness at 17 FPS to maintain real-time simulation. Nonetheless, events are triggered based on the defined contrast threshold and refractory period to ensure that the temporal resolution of events adapts dynamically to scene changes. When the intensity difference at a pixel exceeds an integer multiple of the contrast threshold, a corresponding number of events are generated with evenly distributed timestamps between the 2 frames where that intensity difference was computed. This preserves the temporal granularity of the events, as it produces multiple events during that interval.
Each collected sequence is $45\text{\,}\mathrm{s}$ long, during which the ROV moves and rotates in all degrees of freedom (forward, backward, up, down, and sideways) to capture diverse flow intensities and spatial distributions. These motions are performed at varying speeds up to a maximum velocity of $0.85\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$. A total of 10 sequences were collected. All sequences, along with their corresponding setups, are summarized in Table 1. A sample visualization of the dataset is shown in Figure 3.
@@ 1 Static Constant Reef Down-looking
2 Static Constant Rocky Down-looking
3 Static Varying Reef Down-looking
@@ 4 Static Varying Reef Front-looking
5 Static Varying Rocky Down-looking
@@ 6 Static Varying Rocky Front-looking
7 Dynamic Varying Reef Down-looking
@@ 8 Dynamic Varying Reef Front-looking
9 Dynamic Varying Rocky Down-looking
@@ 10 Dynamic Varying Rocky Front-looking
Table 1: Summary of the generated sequences and their corresponding configurations in the eStonefish-Scenes dataset.
These sequences can be utilized for training neural networks by dividing them into separate event volumes. Each event volume contains a stream of events, along with all grayscale images acquired within the corresponding timestamps. One such example is extracting events that occur between every pair of consecutive grayscale images. The resulting volumes include a grayscale frame at time $t$, another frame at time $t+1$, and all events occurring in between. Based on the 20 FPS refresh rate of the grayscale sensor, we can extract approximately 800 training samples per sequence, providing around 6,400 samples for all sequences combined.
@@ 4 Data Generation Pipeline
The dataset can be easily extended using a modular data generation pipeline, which comprises three separate packages, each responsible for different aspects of data generation. First, we present the Stonefish-SceneGen package. This package is used to generate coral-rich and vibrant scenes by randomly populating the seabed with corals of different types and sizes. It generates a Stonefish-compatible scene file, which can then be used to launch the main pipeline, eStonefish-Scenes. The pipeline is a ROS package with a launch file for Stonefish. The launch file starts the simulator with the generated environment, spawning the ROV and displaying live feeds from all its sensors. The pipeline internally uses the Stonefish-Boids package to randomly spawn schools of fish in the scene. The user can then move the ROV around the scene while recording a rosbag, which can subsequently be converted to the eWiz format using the ConvertStonefish module of eWiz.
Figure 3: Visualisations of event-based, grayscale, and optical flow images for different generated sequences. Event-based images are accumulated over $100\text{\,}\mathrm{ms}$ intervals.
@@ 4.1 Scene Generation
One challenge with the Stonefish simulator is the time-consuming process of manually creating rich and dynamic underwater scenes. Each coral or 3D object must be defined in a scene file, specifying its pose in the world frame, size, and physical attributes. While this approach is feasible for a few objects, it becomes impractical for environments featuring numerous objects. To address this limitation, we introduce Stonefish-SceneGen, an underwater scene generator specifically designed for the Stonefish simulator. This tool streamlines the creation of complex underwater environments, significantly reducing the effort required for scene generation. It consists of a Python script that takes as input a Stonefish scene file, a seafloor surface, and the desired 3D models to spawn. The user specifies the number of coral clusters, along with the number of corals spawned in each cluster. Each coral instance is placed by sampling a random vertex of the seafloor surface. The surface normal at that vertex is then used to determine translation and rotation transformations. To ensure reproducibility, the user can generate the exact same scene by fixing a random seed before running the script. Finally, a scene file containing the path, texture, size, and pose of all spawned 3D models is generated. This scene file can then be used with eStonefish-Scenes to launch the simulation. The scene generator is made openly available at the following link: https://github.com/CIRS-Girona/stonefish-scenegen.
@@ 4.2 Simulating Schools of Fish
To enhance the visual complexity of the underwater scenes with dynamic elements, we incorporated an adapted implementation of the Boids algorithm (boidsReynolds87) for simulating the natural grouping behaviour observed in schools of fish. The algorithm is based on four fundamental rules that govern the behaviour of each fish within the school, collectively influencing group behaviour as a whole.
@@
  • Separation: Each fish receives a repulsion force if it gets too close to another fish in the school, maintaining a minimum distance between individuals.
  • Leader Attraction: One fish is designated as the leader of the school and follows a predefined path through the environment, while the other fish experience an attractive force toward the leader.
@@
To efficiently handle such environmental interactions with static objects, we use OctoMap (octomapHornung13) for discretizing the simulation space, as illustrated in Figure 4. This enables rapid detection of potential collision zones and the assignment of repulsive force vectors to different locations. The flexibility of the library allows for the creation of a custom Octree in which each node stores a repulsive force vector normal to the object surface. When a fish enters an obstacle voxel, a repulsive force is applied, ensuring smooth and natural avoidance behaviour. For dynamic objects such as the ROV, repulsive forces are applied based on the distance between each fish and a defined repulsion radius around the object, allowing for real-time interaction.
@@
Figure 4: Discretised underwater environment represented as an Octree structure.
Additionally, each school of fish follows a randomly generated path created using the Rapidly-exploring Random Trees (RRT) algorithm (rrtSteven98), implemented through the Open Motion Planning Library (OMPL) (omplSucan12). The designated leader follows this path, while the rest of the group moves accordingly based on the leader attraction rule. Each time a new simulation is initiated, both the path and the spawning positions of the fish are randomly generated. However, for reproducibility, users can set a fixed seed to ensure the path and spawning positions remain consistent across runs.
Users can customize the type of fish schools and adjust various behavioural parameters through a YAML configuration file. These include alignment, cohesion, and separation weights, as well as spawning location, model scale, and school size. This functionality is provided as an open-source ROS package, Stonefish-Boids, which builds on the base Stonefish API. The package is available at: https://github.com/CIRS-Girona/stonefish_boids.
@@ 4.3 Data Collection
To record data from the generated scenes, we provide a ROS-based utility, eStonefish-Scenes, which uses the scene file generated by the scene generator and internally calls Stonefish-Boids to simulate schools of fish. It relies on the Stonefish-ROS bridge to facilitate communication with the simulator. The package also includes meshes and textures for the BlueROV2 and a rocky seabed. A predefined scenario combines these assets into a single environment, placing the ROV and the seabed beneath. A Python script is provided to control the robot, utilizing its allocation matrix to enable joystick-based operation. Additionally, an events rendering script is included to visualize the resulting event image in RViz (kamRviz15) as the simulation progresses. Users can customize the sensors mounted on the ROV by modifying the corresponding scene file. Customization options include, but are not limited to, adjusting resolution, field of view, event-based sensor thresholds, as well as adding noise. The record.bash script allows users to record rosbags from the simulation, which can subsequently be converted to the eWiz format using the ConvertStonefish module of eWiz. This data collection utility is available at the following link: https://github.com/CIRS-Girona/estonefish-scenes.
@@ 5 Data Processing
The eStonefish-Scenes dataset was generated with NN training and inference in mind. To streamline this process, we created eWiz, a Python-based library that can be installed with PyPi (pipyInstaller), serving as an all-inclusive framework for event-based data manipulation, processing, visualization, and evaluation. eWiz is designed for seamless integration with libraries such as PyTorch (pytorchLibrary) and Tonic (tonicLibrary), making it suitable for machine learning-oriented pipelines. Additionally, it offers compatibility with existing datasets and supports both event-based and frame-based data. The following subsections present an overview of the most relevant functionalities of eWiz. The library is openly available at the following link: https://github.com/CIRS-Girona/ewiz.
Figure 5: Overview of currently implemented modules in the eWiz library.
@@ 5.1 Data Access and Storage
Due to the high temporal resolution and data throughput of EBCs, event-based data requires substantial storage space. Additionally, slicing data between any two timestamps requires time-consuming search algorithms due to the large volume of events in the sequence. Libraries like NumPy (numpyNumpy) also face high memory demands, as they typically require loading entire sequences into memory before slicing or processing. The eWiz format addresses these challenges with an optimized approach to data management and manipulation. First, we use Blosc (bloscCompression), a high-performance compressor optimized for binary data and effective for compressing numerical arrays. To further reduce data size, the possible values for every event stream was constrained. For instance, pixel coordinates are stored as 16-bit unsigned integers, while polarities are stored as Boolean values. Next, we use the h5py library (h5pyLibrary) to read and write data as HDF5 files (hdf5Format), which support large and complex data structures, and allow slicing without the need to completely load the data into memory. Finally, to reduce loading times, the mappings between timestamps and data indices were pre-computed during the data writing process. This includes binary search to map timestamps (in $ms$) to event, grayscale, and flow indices, as well as direct mappings of event indices to their grayscale and flow counterparts. Additionally, optical flow data is automatically synchronized between specified timestamps by interpolating and accumulating flow displacements as needed.
eWiz provides easy-to-use readers and writers to facilitate this process. The data reader uses the precomputed time mappings to read the desired indices of events based on the start and end timestamps and loads them into memory. It also loads the corresponding grayscale images captured at the beginning and at the end of the event stream, along with the synchronized optical flow. The user can stride over the events using event indices, timestamps, or grayscale image indices. The data reader goes hand in hand with the included data iterators, which allow users to sequentially stride over the data. Such functionality allows the user to obtain event-based data sequentially, without the need for manual indexing.
eWiz also aims to support multiple EBC models to ensure compatibility with publicly available datasets. Several data converters are available for transforming these data types into the eWiz format, including the DAVIS346 (davis346) and Prophesee (prophesee) cameras, along with popular datasets such as MVSEC (mvsecZhu18) and DSEC (dsecGehrig21).